A car radiator has a capacity of 16 quarts and is filled with a 25% Anti-freeze solution. How much must be drained off and replaced with pure Anti-freeze to obtain a 40% Anti-freeze solution?
Answer by ankor@dixie-net.com(12704)

A car radiator has a capacity of 16 quarts and is filled with a 25% Anti-freeze solution. How much must be drained off and replaced with pure Anti-freeze to obtain a 40% Anti-freeze
:
Let x = amt of 25% to be removed
and x = amt of pure antifreeze to be added
:
.25(16-x) + 1.00x = .40(16)
:
4 - .25x + 1x = 6.4
:
-.25x + 1x = 6.4 - 4
:
.75x = 2.4
:
x = 2.4/.75
:
x = 3.2 qts of 25% removed and 3.2 qts of pure antifreeze added
;
;
Check solution, find amt of 25% remaining: 16 - 3.2 = 12.8
.25(12.8) + 3.2 = .4(16)
3.2 + 3.2 = 6.4
;
Did this help, make sense to you?